# Dust evolution during the protostellar collapse: influence on the coupling between the neutral gas and the magnetic field

### v1

@@ 4.1 Impact of the surface energy of the grains $\gamma_{\mathrm{grain}}$
Figure 2 depicts the evolution of the dust mass distribution (first row) and dust number distribution (second row) at different densities during the protostellar collapse, for different values of the grain surface energy. The third row displays the Ohm, ambipolar, Hall resistivities, ion and electron numerical densities evolution with gas density. Likewise for the last row, but with the conductivities and the peak size of the dust distribution. The magnetic resistivities associated to a fixed MRN dust distribution are displayed as a reference in black dashed lines. The impact of the grain surface energy is striking. Indeed, for values of $\gamma_{\mathrm{grain}}=10\ \mathrm{erg\ cm^{-2}}$ and below, fragmentation comes in very early as a replenishment of the population of small grains. It is observed already at $n_{\mathrm{H}}=10^{8}\ \mathrm{cm^{-3}}$. Conversely for larger values, fragmentation only appears at densities higher than $n_{\mathrm{H}}=10^{10}\ \mathrm{cm^{-3}}$, and is even completely absent for $\gamma_{\mathrm{grain}}=190\ \mathrm{erg\ cm^{-2}}$. The peak size of the distribution is lower when decreasing the surface energy: indeed at the end of the collapse, the peak size is of the order of $0.1\ \mathrm{mm}$ for the lowest value of $\gamma_{\mathrm{grain}}$ while it can reach a value of $1\ \mathrm{mm}$ in the case of the largest grain surface energies. In summary, a lower grain surface energy thus entails more fragile grains, which leads to a more rapid and efficient replenishment of the population of small grains along with a reduced peak size of the dust distribution at the end of the collapse.
@@
Such variations in the population of small grains at low densities lead to significant variations in the resistivities. Indeed, for $\gamma_{\mathrm{grain}}=2.9\ \mathrm{erg\ cm^{-2}}$, the immediate replenishment of small grains at all densities gives rise to a subsequent much higher Ohm resistivity throughout the collapse, even at low densities. For $\gamma_{\mathrm{grain}}=10\ \mathrm{erg\ cm^{-2}}$, fragmentation occurs early but slightly later than $\gamma_{\mathrm{grain}}=2.9\ \mathrm{erg\ cm^{-2}}$, and consequently the effects on the Ohm resistivity only emerge later on, as we need the density to sufficiently increase in order for the excess of small grains to collect the electrons. Indeed, a higher number of small grains implies a larger overall cross-section for electron capture by dust grains. This capture reduces the electron density and thus enhances the ohmic resistivity. The ambipolar resistivity $\eta_{\mathrm{AD}}$ is sensitive to the population of small grains as well. In reducing the value of the surface energy of the grains, the small grain re-population results in a higher dust conductivity contribution (which, along with the ions, is the prevailing contribution here, see Appendix A) and thus to a lower ambipolar resistivity $\eta_{\mathrm{AD}}$ at low and intermediate densities. At high densities, the ambipolar resistivity is controlled by the ion density only, and the ambipolar resistivity appears as larger since the excess of small grains reduces the ion density by means of ion-electron recombination at their surface. Finally, the Hall resistivity is very sensitive to the population of small grains and reducing the surface energy induces a re-population of the latter, leading to a transition between negative and positive values that is shifted to higher densities. This transition is even completely suppressed for the lowest value of $\gamma_{\mathrm{grain}}$, which displays a Hall resistivity profile quite similar to that of a fixed MRN distribution (represented as a black dashed line). Indeed, the Hall resistivity is very sensitive to the relative number of electrons and ions. A change in the number of small grains induces a change in the relative number densities of electrons and ions which can easily lead to a change of sign of the Hall resistivity. The more numerous the small grains, the later the transition (change of sign) occurs (see Sect. A for a better description of the Hall conductivity sensitivity to the dust distribution). Depending on the sign of the Hall resistivity, the Hall effect can either promote or reduce the rotation of the collapsing core.
@@
Since the ambipolar resistivity $\eta_{\mathrm{AD}}$ dominates over the Ohm resistivity, we can conclude that more fragile grains (i.e., enhanced fragmentation) leads to a less efficient magnetic diffusion at low and intermediate densities ($n_{\mathrm{H}}\lesssim 10^{10}\ \mathrm{cm^{-3}}$), but more efficient at high density.
@@ 4.2 Exploration for $\gamma_{\mathrm{grain}}=190\ \mathrm{erg\ cm^{-2}}$
We now consider the case of resilient grains for which fragmentation is substantially curtailed.
@@ 4.2.1 Impact of grain-grain erosion
We now explore the impact of grain-grain erosion that can serve as an alternative to fragmentation to repopulate the small grains. We can notice on Fig. 3 the replenishment of the population of very small grains induced by erosion and the limited impact it has on the magnetic resistivities at low density. However, the said impact gets stronger with density as the largest grains grow, allowing for larger grain size ratios and collisional velocities. Recall that according to Eq. (17), the erosion efficiency augments for increasing grain size ratio and grain collisional velocity. In effect, while the model without erosion exhibits a clearly diminishing dust number distribution as the collapse proceeds, the model with erosion maintains a constant population of small grains, with a value dwelling around $\frac{N_{\mathrm{dust}}}{N_{\mathrm{H}}}\leq 10^{-12}$. The rise in erosion efficiency is conspicuous when looking at the mass distribution at the end of the collapse, where not only the tail of the distribution is affected but the distribution as a whole. Consequently, the eroding collisions between the small grains and the largest ones tend to lower the dust peak size at high density. To recapitulate, under standard assumptions (fiducial values for the other parameters) the ambipolar resistivity $\eta_{\mathrm{AD}}$ is only slightly reduced prior to $n_{\mathrm{H}}=10^{11}\ \mathrm{cm^{-3}}$ because the collisions involve grains with three orders of magnitude difference in size at most (collisions between nanometer and micrometer grains). At high density however, the dust distribution spans a much larger range of sizes and thus erosion is much more efficient (the said efficiency increases with the size difference between colliding grains). Consequently, the ambipolar and ohmic resistivities sharply rise, reaching values similar to the MRN case. The Hall resistivity is significantly affected for all densities $n_{\mathrm{H}}>10^{8}\ \mathrm{cm^{-3}}$, where here again, the excess of small grains permitted by erosion processes shifts the change of sign to higher densities. We stress that at high densities, grain-grain erosion significantly reduces the departure from the MRN reference in terms of magnetic resistivities.
4.2.2 Impact of the ambipolar drift intensity $\delta_{\mathrm{AD}}$
@@
We now turn our attention to the impact of the ambipolar drift intensity $\delta_{\mathrm{AD}}$, which controls the amplitude of the collisional velocities induced by ambipolar diffusion (ambipolar drift velocity), as suggested in Eq. (5). In particular, we focus on its combined effect with grain-grain erosion and electrostatic repulsion. Figure 4 depicts the ambipolar resistivity $\eta_{\mathrm{AD}}$ evolution during the protostellar collapse with and without electrostatic repulsion between dust grains, with and without grain-grain erosion for different values of the ambipolar drift intensity. The second row conveys the same information but with a lower grain-grain sticking efficiency ($\epsilon_{\mathrm{strick}}=0.5$). The same plot but for the ohmic resistivity can be found in the Appendix, see Fig. 16.
@@
First of all, by looking at the solid lines of the top left panel, we clearly see that reducing the ambipolar drift between grains alone leads to lower collisional velocities and thus to a much slower coagulation of the small grains. Their number drops less dramatically and consequently ambipolar resistivity $\eta_{\mathrm{AD}}$ is reduced at low density. Then, we turn on electrostatic repulsion (dotted lines) and remark that it has no impact whatsoever on the dust evolution when considering the reference values for the parameters (i.e., $\delta_{\mathrm{AD}}=1$) because while Brownian motion is incapable of generating sufficiently large collisional velocities, the electrostatic barrier is easily overcome by the ambipolar diffusion drift (which allows small grains to be collected by larger grains). However in reducing the ambipolar drift intensity to $\delta_{\mathrm{AD}}=0.1$, we allow the grains to experience collisional velocities lower than the Coulomb velocity threshold, thus inhibiting grain coagulation at low and intermediate densities, as suggested by Fig. 4. Indeed, with this scenario, the difference between the resulting ambipolar resistivity $\eta_{\mathrm{AD}}$ and that associated to an MRN dust distribution is of one order of magnitude at most. Then, as the collapse proceeds and the density increases, the peak size of the dust distribution increases as well and ultimately the largest grains manage to collect the smallest ones via ambipolar drift even for low values of $\delta_{\mathrm{AD}}$, and the three dotted lines end up overlapping. This is due to the fact that the larger the difference in size (or equivalently in Hall factor), the higher the relative collisional velocity induced by ambipolar diffusion (ambipolar drift velocity), see Eq. (6).
@@
In addition, we see on the top right panel that for $\delta_{\mathrm{AD}}=1$, erosion is influential at high density only, as previously highlighted in Sect. 4.2.1. However, for $\delta_{\mathrm{AD}}=10$, while electrostatic repulsion has no impact, grain-grain erosion is rendered very efficient even at low densities due to the large collisional velocities at play. Thus, the significant subsequent replenishment of small grains produces an ambipolar resistivity $\eta_{\mathrm{AD}}$ remarkably lower, close to the MRN case. Those two effects (electrostatic repulsion and erosion) can operate jointly at the same time but respectively allow to reduce the ambipolar resistivity in different circumstances and environments.
@@
Finally, for $\delta_{\mathrm{AD}}=1$, a look at the second row shows that a lower sticking efficiency slows down the coagulation of small grains (see in the Appendix Sect. 17 a more detailed impact of the sticking efficiency). For $\delta_{\mathrm{AD}}=0.1$, the combined effect of electrostatic repulsion and a lower sticking efficiency yields an even more evident reduction of the ambipolar resistivity. For $\delta_{\mathrm{AD}}=10$, the very strong erosion is almost not affected by the reduced sticking efficiency. However, note that the replenishment of small grains induced by erosion at high densities in the case of $\delta_{\mathrm{AD}}=0.1$ and $\delta_{\mathrm{AD}}=1$ is suppressed since half of the collisions lead to bouncing instead of erosion.
@@
In summary, on one hand, if the ambipolar drift is sufficiently weak, erosion has no influence but electrostatic repulsion inhibits small grains depletion as they repel each other instead of sticking. On the other hand, if the ambipolar drift is sufficiently strong, the Coulomb barrier is easily overcome but grain-grain erosion is efficient enough to repopulate the small grains and consequently lower the ambipolar resistivity $\eta_{\mathrm{AD}}$. Note that Appendix D explores the expected values of $\delta_{\mathrm{AD}}$ within a collapsing core.
5 Dust growth: what are the consequences on the magnetic flux ?
@@
Figure 5: Ambipolar resistivity $\eta_{\mathrm{AD}}$ (normalized with the squared initial magnetic field strength) evolution with gas density throughout the protostellar collapse, for different values of the initial magnetic field strength. Solid lines: growing dust distribution ($\gamma_{\mathrm{grain}}=190\ \mathrm{erg\ cm^{-2}}$). Dashed lines : non evolving MRN dust distribution. Dotted red line is the analytical dust free solution.
@@
Figure 7: Freefall timescale (in black) and ambipolar diffusion timescale as a function of gas density throughout the protostellar collapse for different values of the ionization rate induced by cosmic rays. The initial magnetic field chosen corresponds to the critical value (minimum value) above which the small grain depletion leads to an ambipolar diffusion timescale lower than the freefall timescale.
Placing ourselves under standard conditions (reference values for the parameters) but varying the initial magnetic field $B_{0}$ (see Eq. 21) and the ambipolar drift intensity $\delta_{\mathrm{AD}}$, we focus on the evolution of the ambipolar resistivity $\eta_{\mathrm{AD}}$ and on the ambipolar diffusion timescale $t_{\mathrm{AD}}$ (see Eq. 22) during the collapse.
The results of Sect. 4 highlight a tendency for the magnetic resistivities to be significantly different from those one gets with a fixed MRN distribution. In particular, the ambipolar resistivity $\eta_{\mathrm{AD}}$ exhibits high values at low and intermediate densities due to small grain depletion permitted by ambipolar drift. Indeed, since the small grains contribute directly to the perpendicular conductivity at low density (see Sect. A), their removal leads to a sharp increase in the ambipolar resistivity. At high density however, the perpendicular conductivity is controlled by the ionic contribution. In the absence of small grains, ions and electrons do not efficiently recombine and thus the ion density remains high. Therefore at high density, the ambipolar resistivity $\eta_{\mathrm{AD}}$ is much lower relatively to the MRN case. The Ohm resistivity remains lower throughout the collapse, and monotonically increases. Unlike the fixed MRN case, the Hall resistivity systematically switches sign at some point, the corresponding density at which this transition occurs being highly sensitive to the dust distribution.
@@
Placing ourselves under standard assumptions, Fig. 5 depicts the evolution of the ambipolar resistivity $\eta_{\mathrm{AD}}$ during the collapse in the case of resilient grains ($\gamma_{\mathrm{grain}}=190\ \mathrm{erg\ cm^{-2}}$) for different values of the initial magnetic field strength $B_{\mathrm{0}}$. The dashed lines represent the fixed MRN reference case. The resistivity is normalized with $B_{\mathrm{0}}^{2}$ in such a way as to remove the direct dependency in $B^{2}$ and isolate and focus on the influence of the ambipolar diffusion on the resistivity. We wish to see to what extent the $\eta_{\mathrm{AD}}\propto B^{2}$ trend is relevant in the case of an evolving dust distribution. On Fig. 5, we can see that for different values of the initial magnetic field strength, the normalized ambipolar resistivities $\frac{\eta_{\mathrm{AD}}}{B_{\mathrm{0}}^{2}}$ rapidly converge and overlap once the small grains have been removed, which hints at a $B^{2}$ dependency given by the analytical approximation $\eta_{\mathrm{AD}}\propto\frac{B^{2}}{\gamma_{\mathrm{in}}\rho_{\mathrm{i}}\rho_{\mathrm{n}}}$ (Duffin & Pudritz, 2008) with $\gamma_{\mathrm{in}}$ a drag coefficient between ions and neutrals, $\rho_{\mathrm{i}}$ and $\rho_{\mathrm{n}}$ the mass density of the ion and neutral fluids. This analytical expression, which is derived in considering only electrons and ions as charged species, holds true here because the small grains (which contributes to the ambipolar resistivity) are rapidly depleted by the ambipolar diffusion drift. Indeed, this is depicted by the sharp rise in the ambipolar resistivity $\eta_{\mathrm{AD}}$ at the beginning of the collapse, the removal of small grains being even faster for a higher initial magnetic field strength (green curve). Considering a balance between ionization by cosmic-rays and ion/electron recombination, we have
@@
for the numerical ion density when dust is negligible. $\zeta$ is the ionization rate induced by cosmic rays and ${\langle\sigma v\rangle}_{\mathrm{ie}}=2\times 10^{-7}\left(\frac{T}{300\ \mathrm{K}}\right)^{-0.5}\ \mathrm{cm^{-3}\ s^{-1}}$ is the ion/electron recombination rate determined in Marchand et al. (2021). To confirm this $B^{2}$ trend, we use directly the expression of the ambipolar resistivity $\eta_{\mathrm{AD}}$ given by Eq. (3.1.4) to derive a complete expression. The parallel conductivity is dominated by the electron contribution. In absence of dust, the perpendicular one is dominated by the ion and the Hall conductivity is close to zero. Given that $\sigma_{\mathrm{electron}}\gg\sigma_{\mathrm{ion,perp}}$ and $\Gamma_{\mathrm{ion}}\gg 1$ we find that once the small grains have been removed, the ambipolar resistivity scales as

### v2

@@ 4.1 Impact of the surface energy of the grains $\gamma_{\mathrm{grain}}$
Figure 2 depicts the evolution of the dust mass distribution (first row) and dust number distribution (second row) at different densities during the protostellar collapse, for different values of the grain surface energy. The third row displays the Ohm, ambipolar, Hall resistivities, ion and electron numerical densities evolution with gas density. Likewise for the last row, but with the conductivities and the peak size of the dust distribution. The magnetic resistivities associated with a fixed MRN dust distribution are displayed as a reference in black dashed lines. The impact of the grain surface energy is striking. Indeed, for values of $\gamma_{\mathrm{grain}}=10\ \mathrm{erg\ cm^{-2}}$ and below, fragmentation comes in very early as a replenishment of the population of small grains. It is observed already at $n_{\mathrm{H}}=10^{8}\ \mathrm{cm^{-3}}$. Conversely for larger values, fragmentation only appears at densities higher than $n_{\mathrm{H}}=10^{10}\ \mathrm{cm^{-3}}$, and is even completely absent for $\gamma_{\mathrm{grain}}=190\ \mathrm{erg\ cm^{-2}}$. The peak size of the distribution is lower when decreasing the surface energy: indeed at the end of the collapse, the peak size is of the order of $0.1\ \mathrm{mm}$ for the lowest value of $\gamma_{\mathrm{grain}}$ while it can reach a value of $1\ \mathrm{mm}$ in the case of the largest grain surface energies. In summary, a lower grain surface energy thus entails more fragile grains, which leads to a more rapid and efficient replenishment of the population of small grains along with a reduced peak size of the dust distribution at the end of the collapse.
@@
Such variations in the population of small grains at low densities lead to significant variations in the resistivities. Indeed, for $\gamma_{\mathrm{grain}}=2.9\ \mathrm{erg\ cm^{-2}}$, the immediate replenishment of small grains at all densities gives rise to a subsequent much higher Ohm resistivity throughout the collapse, even at low densities. For $\gamma_{\mathrm{grain}}=10\ \mathrm{erg\ cm^{-2}}$, fragmentation occurs early but slightly later than $\gamma_{\mathrm{grain}}=2.9\ \mathrm{erg\ cm^{-2}}$, and consequently the effects on the Ohm resistivity only emerge later on, as we need the density to sufficiently increase in order for the excess of small grains to collect the electrons. Indeed, a higher number of small grains implies a larger overall cross-section for electron capture by dust grains. This capture reduces the electron density and thus enhances the ohmic resistivity. The ambipolar resistivity $\eta_{\mathrm{AD}}$ is sensitive to the population of small grains as well. In reducing the value of the surface energy of the grains, the small grain re-population results in a higher dust conductivity contribution (which, along with the ions, is the prevailing contribution here; see Appendix A) and thus to a lower ambipolar resistivity $\eta_{\mathrm{AD}}$ at low and intermediate densities. At high densities, the ambipolar resistivity is controlled by the ion density only, and the ambipolar resistivity appears as larger since the excess of small grains reduces the ion density by means of ion-electron recombination at their surface. Finally, the Hall resistivity is very sensitive to the population of small grains and reducing the surface energy induces a re-population of the latter, leading to a transition between negative and positive values that is shifted to higher densities. This transition is even completely suppressed for the lowest value of $\gamma_{\mathrm{grain}}$, which displays a Hall resistivity profile quite similar to that of a fixed MRN distribution (represented as a black dashed line). Indeed, the Hall resistivity is very sensitive to the relative number of electrons and ions. A change in the number of small grains induces a change in the relative number densities of electrons and ions which can easily lead to a change of sign of the Hall resistivity. The more numerous the small grains, the later the transition (change of sign) occurs (see Sect. A for a better description of the Hall conductivity sensitivity to the dust distribution). Depending on the sign of the Hall resistivity, the Hall effect can either promote or reduce the rotation of the collapsing core.
@@
Since the ambipolar resistivity $\eta_{\mathrm{AD}}$ dominates over the Ohm resistivity, we can conclude that more fragile grains (i.e. enhanced fragmentation) leads to a less efficient magnetic diffusion at low and intermediate densities ($n_{\mathrm{H}}\lesssim 10^{10}\ \mathrm{cm^{-3}}$), but more efficient at high density.
@@ 4.2 Exploration for $\gamma_{\mathrm{grain}}=190\ \mathrm{erg\ cm^{-2}}$
We go on to consider the case of resilient grains for which fragmentation is substantially curtailed.
@@ 4.2.1 Impact of grain-grain erosion
Here, we explore the impact of grain-grain erosion that can serve as an alternative to fragmentation to repopulate the small grains. We note on Fig. 3 the replenishment of the population of very small grains induced by erosion and the limited impact it has on the magnetic resistivities at low density. However, the said impact gets stronger with density as the largest grains grow, allowing for larger grain size ratios and collisional velocities. We recall that according to Eq. (17), the erosion efficiency augments for increasing grain size ratio and grain collisional velocity. In effect, while the model without erosion exhibits a clearly diminishing dust number distribution as the collapse proceeds, the model with erosion maintains a constant population of small grains, with a value hovering around $\frac{N_{\mathrm{dust}}}{N_{\mathrm{H}}}\leq 10^{-12}$. The rise in erosion efficiency is conspicuous when looking at the mass distribution at the end of the collapse, where not only the tail of the distribution is affected but the distribution as a whole. Consequently, the eroding collisions between the small grains and the largest ones tend to lower the dust peak size at high density. To recapitulate, under standard assumptions (fiducial values for the other parameters) the ambipolar resistivity, $\eta_{\mathrm{AD}}$, is only slightly reduced prior to $n_{\mathrm{H}}=10^{11}\ \mathrm{cm^{-3}}$ because the collisions involve grains with three orders of magnitude difference in size at most (collisions between nanometer and micrometer grains). At high density however, the dust distribution spans a much larger range of sizes and thus erosion is much more efficient (the said efficiency increases with the size difference between colliding grains). Consequently, the ambipolar and ohmic resistivities sharply rise, reaching values similar to the MRN case. The Hall resistivity is significantly affected for all densities $n_{\mathrm{H}}>10^{8}\ \mathrm{cm^{-3}}$. Again, in this case, the excess of small grains enabled by erosion processes shifts the change of sign to higher densities. We stress that at high densities, grain-grain erosion significantly reduces the departure from the MRN reference in terms of magnetic resistivities.
4.2.2 Impact of the ambipolar drift intensity, $\delta_{\mathrm{AD}}$
@@
We now turn our attention to the impact of the ambipolar drift intensity $\delta_{\mathrm{AD}}$, which controls the amplitude of the collisional velocities induced by ambipolar diffusion (ambipolar drift velocity), as suggested in Eq. (5). In particular, we focus on its combined effect with grain-grain erosion and electrostatic repulsion. Figure 4 depicts the ambipolar resistivity $\eta_{\mathrm{AD}}$ evolution during the protostellar collapse with and without electrostatic repulsion between dust grains, with and without grain-grain erosion for different values of the ambipolar drift intensity. The second row conveys the same information but with a lower grain-grain sticking efficiency ($\epsilon_{\mathrm{strick}}=0.5$). The same plot but for the ohmic resistivity can be found in the appendix (see Fig. 16).
@@
First of all, by looking at the solid lines of the top left panel, we clearly see that reducing the ambipolar drift between grains alone leads to lower collisional velocities and thus to a much slower coagulation of the small grains. Their number drops less dramatically and consequently ambipolar resistivity $\eta_{\mathrm{AD}}$ is reduced at low density. Then, we turn on electrostatic repulsion (dotted lines) and remark that it has no impact whatsoever on the dust evolution when considering the reference values for the parameters (i.e. $\delta_{\mathrm{AD}}=1$) because while Brownian motion is incapable of generating sufficiently large collisional velocities, the electrostatic barrier is easily overcome by the ambipolar diffusion drift (which allows for small grains to be collected by larger grains). However in reducing the ambipolar drift intensity to $\delta_{\mathrm{AD}}=0.1$, we allowed the grains to experience collisional velocities lower than the Coulomb velocity threshold, thus inhibiting grain coagulation at low and intermediate densities, as suggested by Fig. 4. Indeed, with this scenario, the difference between the resulting ambipolar resistivity $\eta_{\mathrm{AD}}$ and that associated with an MRN dust distribution is of one order of magnitude at most. Then, as the collapse proceeds and the density increases, the peak size of the dust distribution increases as well and ultimately the largest grains manage to collect the smallest ones via ambipolar drift even for low values of $\delta_{\mathrm{AD}}$, and the three dotted lines end up overlapping. This is due to the fact that the larger the difference in size (or equivalently in Hall factor), the higher the relative collisional velocity induced by ambipolar diffusion (ambipolar drift velocity), see Eq. (6).
@@
In addition, we see on the top right panel that for $\delta_{\mathrm{AD}}=1$, erosion is influential at high density only, as previously highlighted in Sect. 4.2.1. However, for $\delta_{\mathrm{AD}}=10$, while electrostatic repulsion has no impact, grain-grain erosion is rendered very efficient even at low densities due to the large collisional velocities at play. Thus, the significant subsequent replenishment of small grains produces an ambipolar resistivity $\eta_{\mathrm{AD}}$ remarkably lower, close to the MRN case. Those two effects (electrostatic repulsion and erosion) can operate jointly at the same time but respectively allow us to reduce the ambipolar resistivity in different circumstances and environments.
@@
Finally, for $\delta_{\mathrm{AD}}=1$, a look at the second row shows that a lower sticking efficiency slows down the coagulation of small grains (see Appendix 17 for a more detailed impact of the sticking efficiency). For $\delta_{\mathrm{AD}}=0.1$, the combined effect of electrostatic repulsion and a lower sticking efficiency yields an even more evident reduction of the ambipolar resistivity. For $\delta_{\mathrm{AD}}=10$, the very strong erosion is almost not affected by the reduced sticking efficiency. However, note that the replenishment of small grains induced by erosion at high densities in the case of $\delta_{\mathrm{AD}}=0.1$ and $\delta_{\mathrm{AD}}=1$ is suppressed since half of the collisions lead to bouncing instead of erosion.
@@
In summary, on one hand, if the ambipolar drift is sufficiently weak, erosion has no influence but electrostatic repulsion inhibits small grains depletion as they repel each other instead of sticking. On the other hand, if the ambipolar drift is sufficiently strong, the Coulomb barrier is easily overcome but grain-grain erosion is efficient enough to repopulate the small grains and consequently lower the ambipolar resistivity $\eta_{\mathrm{AD}}$. We note that Appendix D explores the expected values of $\delta_{\mathrm{AD}}$ within a collapsing core.
5 Dust growth: Potential consequences for the magnetic flux
@@
Figure 5: Ambipolar resistivity, $\eta_{\mathrm{AD}}$, (normalized with the squared initial magnetic field strength) evolution with gas density throughout the protostellar collapse, for different values of the initial magnetic field strength. Solid lines: growing dust distribution ($\gamma_{\mathrm{grain}}=190\ \mathrm{erg\ cm^{-2}}$). Dashed lines : non evolving MRN dust distribution. Dotted red line is the analytical dust-free solution.
@@
Figure 7: Freefall timescale (in black) and ambipolar diffusion timescale as a function of gas density throughout the protostellar collapse for different values of the ionisation rate induced by cosmic rays. The initial magnetic field chosen corresponds to the critical value (minimum value) above which the small grain depletion leads to an ambipolar diffusion timescale lower than the freefall timescale.
Assuming standard conditions (reference values for the parameters), but varying the initial magnetic field $B_{0}$ (see Eq. 21) and the ambipolar drift intensity $\delta_{\mathrm{AD}}$, we can focus on the evolution of the ambipolar resistivity, $\eta_{\mathrm{AD}}$, and on the ambipolar diffusion timescale, $t_{\mathrm{AD}}$, (see Eq. 22) during the collapse.
The results of Sect. 4 highlight a tendency for the magnetic resistivities to be significantly different from those one gets with a fixed MRN distribution. In particular, the ambipolar resistivity $\eta_{\mathrm{AD}}$ exhibits high values at low and intermediate densities due to small grain depletion enabled by ambipolar drift. Indeed, since the small grains contribute directly to the perpendicular conductivity at low density (see Sect. A), their removal leads to a sharp increase in the ambipolar resistivity. At high density however, the perpendicular conductivity is controlled by the ionic contribution. In the absence of small grains, ions, and electrons do not efficiently recombine and thus the ion density remains high. Therefore at high density, the ambipolar resistivity, $\eta_{\mathrm{AD}}$, is much lower with respect to the MRN case. The Ohm resistivity remains lower throughout the collapse, and monotonically increases. Unlike the fixed MRN case, the Hall resistivity systematically switches sign at some point, the corresponding density at which this transition occurs being highly sensitive to the dust distribution.
@@
Placing ourselves under standard assumptions, Fig. 5 depicts the evolution of the ambipolar resistivity $\eta_{\mathrm{AD}}$ during the collapse in the case of resilient grains ($\gamma_{\mathrm{grain}}=190\ \mathrm{erg\ cm^{-2}}$) for different values of the initial magnetic field strength $B_{\mathrm{0}}$. The dashed lines represent the fixed MRN reference case. The resistivity is normalized with $B_{\mathrm{0}}^{2}$ in such a way as to remove the direct dependency in $B^{2}$ and isolate and focus on the influence of the ambipolar diffusion on the resistivity. We wish to see to what extent the $\eta_{\mathrm{AD}}\propto B^{2}$ trend is relevant in the case of an evolving dust distribution. On Fig. 5, we can see that for different values of the initial magnetic field strength, the normalized ambipolar resistivities $\frac{\eta_{\mathrm{AD}}}{B_{\mathrm{0}}^{2}}$ rapidly converge and overlap once the small grains have been removed, which hints at a $B^{2}$ dependency given by the analytical approximation, $\eta_{\mathrm{AD}}\propto\frac{B^{2}}{\gamma_{\mathrm{in}}\rho_{\mathrm{i}}\rho_{\mathrm{n}}}$ (Duffin & Pudritz, 2008), with $\gamma_{\mathrm{in}}$ as the drag coefficient between ions and neutrals, then $\rho_{\mathrm{i}}$ and $\rho_{\mathrm{n}}$ as the mass density of the ion and neutral fluids. This analytical expression, which is derived in considering only electrons and ions as charged species, holds true here because the small grains (which contributes to the ambipolar resistivity) are rapidly depleted by the ambipolar diffusion drift. Indeed, this is depicted by the sharp rise in the ambipolar resistivity, $\eta_{\mathrm{AD}}$, at the beginning of the collapse, the removal of small grains being even faster for a higher initial magnetic field strength (green curve). Considering a balance between ionisation by cosmic-rays and ion/electron recombination, we have
@@
for the numerical ion density when dust is negligible. Here, $\zeta$ is the ionisation rate induced by cosmic rays and ${\langle\sigma v\rangle}_{\mathrm{ie}}=2\times 10^{-7}\left(\frac{T}{300\ \mathrm{K}}\right)^{-0.5}\ \mathrm{cm^{-3}\ s^{-1}}$ is the ion/electron recombination rate determined in Marchand et al. (2021). To confirm this $B^{2}$ trend, we use directly the expression of the ambipolar resistivity $\eta_{\mathrm{AD}}$ given by Eq. (3.1.4) to derive a complete expression. The parallel conductivity is dominated by the electron contribution. In absence of dust, the perpendicular one is dominated by the ion and the Hall conductivity is close to zero. Given that $\sigma_{\mathrm{electron}}\gg\sigma_{\mathrm{ion,perp}}$ and $\Gamma_{\mathrm{ion}}\gg 1$ we find that once the small grains have been removed, the ambipolar resistivity scales as